40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Fri Nov 6 08:14:58 2020, Aidan, Update, 2um Photodiodes, Photodiode testing recovery status 
    Reply  Fri Nov 6 12:45:21 2020, Aidan, Update, 2um Photodiodes, Photodiode testing recovery status MEDM.pngPD_sweep.png
Message ID: 2514     Entry time: Fri Nov 6 12:45:21 2020     In reply to: 2513
Author: Aidan 
Type: Update 
Category: 2um Photodiodes 
Subject: Photodiode testing recovery status 

Embellished Chris's PD MEDM screen a bit to illustrate controls in a diagram. The representation of the RELAY SWITCH between the Keithley and the SR560 is a bit off - I think the transimpedance amplifier is switched out as well.

Also - Keithley bright PD sweep output is attached.

Quote:

Quick update, more detailed update to follow.

  • Laser is working
  • Photodiode sweep with the Keithley shows a sensible dark Current v Voltage plot (when laser is off) - indicating that PD wiring is still intact
  • Laser was aligned onto photodiode (although it took a while aligning to find the signal)
  • Ran a sweep with the Keithley and the laser on - saw another sensible bright Current v Voltage plot (more current than in the dark case)
  • DAQ control still works
  • DAC output is directly providing (unfiltered) 200mV bias

Still to do:

  • Get the SR785 plugged back in
  • Get an SR560 inserted between DAC output and PD bias to low pass filter
  • Investigate why the laser current set point is so noisy
  • Sort out RTD situation inside the chamber
  • Miscellaneous stuff

 

Attachment 1: MEDM.png  27 kB  | Hide | Hide all
MEDM.png
Attachment 2: PD_sweep.png  26 kB  | Hide | Hide all
PD_sweep.png
ELOG V3.1.3-